5x – 2y = 4
how do you do slope

Found 2 solutions by stanbon, roseo:
Answer by stanbon(75887)

5x – 2y = 4
how do you do slope
-------------------------
Solve for "Y":
5x = 2y + 4
2y = 5x-4
y = (5/2)x -2
slope = (5/2)
=================
Cheers,
Stan H.

Answer by roseo(33)

5x-2y = 4
Put equation in the slope intercept form (y = mx +b) by solving the equation for y.
-2y = -5x +4
y = (5/2)x - 2
Because m is the slope in y=mx+b then 5/2 is the slope.
An alternative method is to find two points that are one the line.
Say pick x=0 then y = -2 (Substitute 0 in for x and then solve for y)
Say pick x=2 then y=3 (Substitute 2 in for x and then solve for y)
Now you have the two points (0,-2) and (2,3) and use the slope formula
3 - (-2)
________
2-0
=5/2